OpenStudy (anonymous):

can anyone help me with this problem find vertex and equation, line of symmetry and graph the function. f(x)= 1/3 x^2

OpenStudy (anonymous):

f(x)=\[\frac{ x ^{2} }{ 3 }\]

OpenStudy (anonymous):

I might be reading it wrong I am sorry

OpenStudy (anonymous):

the parabola or quadratic equation is f(x)=ax^2 +bx +c, vertex x=-b/2a if you have a formula of f(x)= 1/3 x^2+bx +c --> here b=0 and c=0 vertex x=-b/2a x=-0/2(1/3)=0 sub this to f(x)= y=1/3 x^2=0 then what is the vertex V(x,y)=___? is it V(0,0) yes or no? for graphing use values of x=0,+-1,+-2 +-3 etc..etc....:D good luck now

OpenStudy (anonymous):

You had it right I think I was confused, but you really helped me understand the question...

OpenStudy (anonymous):

ok good,,, good luck now and have fun ....:D

OpenStudy (anonymous):

I will once someone looks at my problems lol..... One person is looking I just want to ensure I am on the right track....

OpenStudy (anonymous):

ok where is the prob? i may be able to help a little bit :D

OpenStudy (anonymous):

How would I show you?

OpenStudy (anonymous):

If anyone can help me check my work I would be very thankful. So it can be more understandable I attached my work in a word file. Thank you so very much I just want to ensure I am on the right track.

OpenStudy (anonymous):

Thats the question

OpenStudy (anonymous):

hmm go to your prob site and notify my name there then i will click it and be there :D

OpenStudy (anonymous):

like this hi andjie

OpenStudy (anonymous):

I say your name

OpenStudy (anonymous):

yes copy and paste my name there then ill just click on it

OpenStudy (anonymous):

ok I did that

OpenStudy (anonymous):

hmm i didnt have a notification,, did you highlight and copy then paste my name there?

OpenStudy (anonymous):

Paste it where exactly lol

OpenStudy (anonymous):

in the type your reply area, because I did that

OpenStudy (anonymous):

on where you posted your problem

OpenStudy (anonymous):

ok go back to where you posted your problem then paste my name there

OpenStudy (anonymous):

I posted your name in the question

OpenStudy (anonymous):

So confusing

OpenStudy (anonymous):

hmm i dont know why i didnt get a notification?

OpenStudy (anonymous):

is it still open? why dont you close it hen repost them new

OpenStudy (anonymous):

ok